one vertex of a rectangle is at (1,2) and the point of intersection of the diagnols is at (5,4). what are the coordnates of the other vertices of the rectangle if two sides of the rectangle are parallel to the x-axis
Answer by ewatrrr(24785)


Hi,
Graphing demonstates: other vetices are (1,6),(9,2) and(9,6)
One could also in general use the slope of the diagonal containing:
(1,2) and (5,4) = 4-2 / 5-1 = 2/4 to find the other vertices